# Tazza: Shuffling Neural Network Parameters for Secure and Private Federated Learning

### v1

@@ Abstract.
Federated learning enables decentralized model training without sharing raw data, preserving data privacy. However, its vulnerability towards critical security threats, such as gradient inversion and model poisoning by malicious clients, remain unresolved. Existing solutions often address these issues separately, sacrificing either system robustness or model accuracy. This work introduces Tazza, a secure and efficient federated learning framework that simultaneously addresses both challenges. By leveraging the permutation equivariance and invariance properties of neural networks via weight shuffling and shuffled model validation, Tazza enhances resilience against diverse poisoning attacks, while ensuring data confidentiality and high model accuracy. Comprehensive evaluations on various datasets and embedded platforms show that Tazza achieves robust defense with up to 6.7$\times$ improved computational efficiency compared to alternative schemes, without compromising performance.
@@ 1. Introduction
Federated learning is a transformative paradigm that facilitates training machine learning models in decentralized environments while preserving the privacy of data stored on distributed devices (McMahan et al., 2017). In federated learning, clients perform local model training over several epochs and share model updates, rather than raw data, with a central server. The server then aggregates these updates across clients over multiple rounds, collaboratively building a robust global model. This decentralized approach is particularly promising for mobile and embedded computing systems (Shin et al., 2024; Alam et al., 2023), enabling them to leverage vast amounts of locally collected data while delivering personalized models tailored to individual users (Sani et al., 2024).
Despite its decentralized and privacy-preserving design, federated learning still faces significant privacy and security threats. Unlike centralized training, where data resides on a single server, federated learning relies on distributed, heterogeneous client data, leaving it vulnerable to malicious model updates that compromise system integrity (Fang et al., 2020). Furthermore, honest-but-curious servers can analyze model updates to infer private client data, breaching confidentiality. These concerns are particularly critical in mobile and embedded systems, where sensitive data is abundant, and the software environment is vulnerable due to the ease of installing third-party apps. These threats are categorized into two primary attack vectors: integrity attacks, in which malicious clients undermine model robustness, and confidentiality attacks, where curious servers endanger data privacy.
@@
We emphasize that ensuring both integrity and confidentiality is critical in federated learning, particularly for mobile and embedded systems. With large-scale participation from distributed and often anonymous clients, it is unrealistic to assume that all participants act in good faith. Robust mechanisms are needed to defend against malicious clients that compromise the integrity of the system. Equally important is safeguarding sensitive data collected from mobile and embedded devices, such as health information (Park et al., 2020), daily activities (Park et al., 2021; Ouyang et al., 2024; Fu et al., 2024), or environmental metrics (Xie et al., 2023; Sun et al., 2020) to maintain user trust and meet privacy regulations.
However, addressing integrity and confidentiality simultaneously poses unique challenges. Confidentiality defenses, such as noise addition, obscure model updates to protect privacy but can inadvertently hinder the detection of malicious patterns for backdoor attacks or label poisoning. At the same time, integrity defenses identify and mitigate such threats but may conflict with confidentiality measures, complicating the design of robust systems that target both objectives.
This challenge is amplified by the non-i.i.d. nature of real-world data. The diversity of benign model updates makes it difficult to differentiate legitimate updates from malicious ones, especially when compounded by the inaccuracies introduced by confidentiality defenses. These trade-offs create significant barriers to ensuring security and privacy without degrading global model performance.
Despite the critical importance of these issues, few approaches simultaneously address integrity and confidentiality threats while preserving model performance. Tackling this intertwined challenge is imperative for real-world federated learning applications, where system reliability and user trust depend on effective, balanced defense mechanisms.
@@
Leveraging these properties, we propose Tazza, a novel approach that simultaneously ensures confidentiality and integrity in federated learning. Specifically, by shuffling neural network model parameters, Tazza utilizes the permutation equivariance property to obscure sensitive data, enhancing confidentiality. Meanwhile, permutation invariance ensures that the shuffled parameters produce consistent computational outcomes, thereby maintaining integrity. This approach effectively demonstrates how permutation-based techniques can address fundamental security issues in federated learning without sacrificing model performance.
Nevertheless, properly shuffling model parameters while preserving desired properties is challenging, as it requires a deep understanding of neural network operations. While many architectures exhibit permutation equivariant and invariant properties, the mechanisms enabling these properties vary across models. To address this, we present tailored algorithms for representative architectures, offering a systematic approach to harness these properties. Our methods demonstrate the practical potential of permutation-based techniques in federated learning.
Building on the aforementioned properties, Tazza introduces a novel weight shuffling module, designed to obscure the original model by rearranging its weights while maintaining permutation equivariant and invariant properties. This ensures that the shuffled model remains valid, enabling us to propose a shuffled validation process for integrity checks, confirming that model outputs remain accurate despite weight shuffling. Finally, we leverage the similarity metrics obtained from the shuffled validation process and cluster the models to distinguish and isolate malicious updates from impacting benign models.
By integrating these components, Tazza emerges as a secure, privacy-preserving, and accurate federated learning framework that effectively addresses both integrity and confidentiality challenges. Along with a strong mathematical foundation leveraging permutation properties, we demonstrate its practical efficacy through extensive evaluations on diverse datasets and embedded platforms. Specifically, Tazza maintains high model accuracy while mitigating both integrity and confidentiality attacks, with a computational speedup of up to 6.7$\times$ compared to alternative defense schemes. Our key contributions are summarized as follows:
@@ 2. Background and Related Works
2.1. Security Threats in Federated Learning
We begin by discussing background information on the two core security threats in federated learning: integrity and confidentiality attacks.
To mitigate integrity attacks, prior research has proposed secure model aggregation techniques leveraging statistical priors or similarity-based approaches. Methods like Median (Xie et al., 2018) and Trimmed-Mean (Yin et al., 2018) aggregation exclude malicious updates by focusing on statistical measures rather than naive averaging, assuming that malicious updates are statistical outliers. Alternatively, similarity-based techniques such as Krum (Blanchard et al., 2017) use Euclidean distance to select updates with minimal deviation across models, while FLTrust (Cao et al., 2020) calculates trust scores based on cosine similarity with a verified, clean model stored at the server. Despite their strengths, these methods often discard valuable information in the updates, leading to suboptimal global model performance.
To counter confidentiality attacks, widely adopted approaches include applying differential privacy (Abadi et al., 2016), which adds controlled noise to model updates, and gradient pruning (Zhu et al., 2019), which suppresses sensitive information in gradients during local training. However, these defense mechanisms often degrade model performance, as they manipulate original updates in ways that hinder convergence and accuracy (Huang et al., 2021).
2.2. Permutation Equivariance/Invariance
@@
Permutation Equivariance refers to the property where the order of a function’s output aligns with the order of its input. Figure 2 (a) illustrates this concept using matrix multiplication, where $R$ represents a column-reversed matrix. Specifically, the figure demonstrates that when the column order of the input matrix $X$ is reversed ($X\rightarrow X^{R}$), the resulting matrix multiplication outputs, $WX$ (top) and $WX^{R}$ (bottom), differ only in the column order, which corresponds to the input’s order.
Formally, if a function $f(\cdot)$ satisfies permutation equivariance, for a given input $X=\{x_{1},x_{2},\dots,x_{N}\}$ and a permutation function $\pi(i):\{1,2,\ldots,N\}\to\{1,2,\ldots,N\}$, where $\pi(i)$ is a bijective mapping, the lemma $\pi(f(X))=f(\pi(X))$ holds. Applying this to the case of matrix multiplication, the function $f(X)=WX$ satisfies permutation equivariance with respect to the column order. Consequently, $WX^{R}$ is identical to $(WX)^{R}$.
Permutation Invariance describes a function that produces identical results regardless of the order of its input. For example, a function that computes the average of the elements in a given array is permutation invariant because the arrangement of elements does not influence the outcome. Figure 2 (b) illustrates this concept using the average operation in the context of matrix multiplication, where $n$ represents the number of rows and $R$ denotes a row-permuted version of the input. As shown, despite reordering the rows, the final result remains unchanged (i.e., $\mathbb{E}[X]=\mathbb{E}[X^{R}]$). Without loss of generality, if a function $f(\cdot)$ satisfies permutation invariance, then for a given input $X$ and a permutation function $\pi(i)$, the lemma $\pi(f(X))=f(\pi(X))$ holds.
It is important to note that matrix multiplication, in general, is not a permutation invariant operation. Specifically, if the weights used in the multiplication are not uniform (e.g., $\frac{1}{n}$ for all elements, as in the averaging operation shown in Figure 2 (b)), the result will differ depending on the order of the rows. This is because such weights can be interpreted as a weighted sum, being sensitive to the input arrangement.
@@ 3. Feasibility Study and Core Idea
Figure 3. (a) Bottom-4 (top green box) and Top-4 (bottom red box) PSNR samples from the MNIST dataset. (b) PSNR distribution between the original and shuffled images from the MNIST dataset (using 60K images).
Feasibility Study. To evaluate the feasibility of leveraging shuffling methods for preserving data privacy, we measured the Peak Signal-to-Noise Ratio (PSNR) between original and shuffled samples in the MNIST dataset. For shuffling, we randomly rearranged the pixels in the input data following a consistent and fixed order applied uniformly across all samples. Figure 3 presents sample original/shuffled images and the distribution of measured PSNR values between 60,000 original and shuffled images. A high PSNR indicates structural similarity between two images, whereas a low PSNR suggests significant differences. As Figure 3 (a) shows, the shuffled samples do not retain the semantic information of the original images. Moreover, Figure 3 (b) shows that most samples exhibit low PSNR in the range of 6–8 dB. Even for the top-4 PSNR samples, the shuffled images remain semantically uninterpretable to human observers, confirming the effectiveness of shuffling in preserving data privacy.
@@
Next, we trained a 3-layer MLP model for 20 epochs using two dataset configurations: one with original data and the other with shuffled data, as exemplified in Figure 3 (a). For the shuffling process, the test data and training data were shuffled using the same order. Figure 4 presents the average validation accuracy for 10 different random seeds. As the plots show, the model achieves (nearly) identical accuracy regardless of whether the input data were shuffled, demonstrating the neural network’s semi-permutation invariance property. This result suggests that data shuffling can preserve computation consistency in both training and inference, while effectively addressing privacy concerns.
@@ 4. Threat Model
Goals and Capability First, an integrity attacker refers to a malicious client that aims to compromise the performance of the global model by injecting harmful information through tampered model updates. Integrity attackers can launch model or data poisoning attacks, manipulating the local training process by altering labels, training data, or model weights. Second, we consider an honest-but-curious server as a confidentiality attacker, which seeks to infer a client’s training data by exploiting the received information. Given that clients transmit their model updates to the server for aggregation, the attacker has access to the updates explicitly shared by the clients. Furthermore, there can be additional shared information, which can be any type of data that the clients are required to transmit for the global model training process (i.e., amount of training data used).
Assumptions For the integrity attack, we assume that the attacker has no control over its device beyond the local training process. Specifically, the attacker cannot interfere with the weight aggregation, distribution, or client selection processes carried out by the federated learning server, nor can it access information from other benign clients, such as their data distribution, model weights, or updates. Finally, we assume that the proportion of attackers cannot exceed 50% of the total number of clients (Fang et al., 2020), all being common assumptions in prior work. In the case of a confidentiality attacker, the attacker cannot interfere with or access the local operations performed on clients, nor can it eavesdrop on network communications. Lastly, we note that both types of attackers make every effort to achieve their objectives and circumvent defense mechanisms, while the victim systems operate cooperatively to prevent these attacks (Kabir et al., 2024).

### v2

@@ Abstract.
Federated learning is an innovative computing paradigm that enables decentralized model training without sharing raw data from individual user/client devices, thereby preserving their data privacy. However, federated learning systems still face significant security challenges, from attacks including data leakage via gradient inversion and model poisoning from malicious clients. Existing solutions often address these issues in isolation, which compromises either system robustness or model accuracy. In this paper, we present Tazza, a secure and efficient federated learning framework designed to tackle both challenges simultaneously. By leveraging the permutation equivariance and invariance property inherent in modern neural network architectures, Tazza enhances resilience against diverse poisoning attacks, including noise injection and backdoor attacks, while ensuring data confidentiality and maintaining high model accuracy. Evaluations using various datasets and embedded computing platforms show that Tazza achieves robust defense with up to 6.7$\times$ improved computational efficiency compared to alternative schemes, without compromising the federated learning model performance.
@@ 1. Introduction
Federated learning is a transformative paradigm that facilitates training machine learning models in decentralized environments while preserving the privacy of data stored on distributed mobile and embedded devices (McMahan et al., 2017) and enhancing personalized user experiences in ubiquitous computing applications. In federated learning, clients perform local model training over several epochs and share model updates, rather than raw data, with a central server. The server then aggregates these updates across clients over multiple rounds, collaboratively building a robust global model. This decentralized approach is particularly promising for mobile and embedded computing systems (Shin et al., 2024; Alam et al., 2023), given that it enables them to leverage vast amounts of (distributed) locally collected data for maintaining a high-performing personalized model while alleviating many of the privacy and network overhead concerns that arise from sharing personal data/information (Sani et al., 2024).
However, despite its decentralized and privacy-preserving nature stemming from the fact that local information of individual users/devices is not shared directly with the server, federated learning still faces significant privacy and security threats. Unlike centralized training, where data resides on a single server, federated learning relies on distributed, heterogeneous user/client data, leaving it vulnerable to malicious model updates that compromise system integrity (Fang et al., 2020; Elhattab et al., 2023). Furthermore, honest-but-curious servers can attempt to analyze model updates to infer private user data, breaching confidentiality (Elhattab et al., 2024; Feng et al., 2020). These concerns are particularly critical in ubiquitous computing scenarios, such as daily activity and personal health monitoring, where sensitive data is abundant, and the software environment is vulnerable due to the ease of installing third-party apps (Lee et al., 2025, 2023). These threats are categorized into two primary attack vectors: integrity attacks, in which malicious clients undermine model robustness, and confidentiality attacks, where curious servers endanger data privacy.
@@
We emphasize that ensuring both integrity and confidentiality is critical in federated learning, particularly for mobile and embedded systems. With large-scale participation from distributed and often anonymous clients, it is unrealistic to assume that all participants act in good faith. Robust mechanisms are needed to defend against malicious clients that compromise the integrity of the system. Equally important is safeguarding sensitive data collected from mobile and embedded devices, such as health information (Park et al., 2020; Ouyang et al., 2024), daily activities (Park et al., 2021; Fu et al., 2024; Yun et al., 2024), or environmental metrics (Xie et al., 2023; Sun et al., 2020) to maintain user trust and meet privacy regulations.
However, simultaneously addressing integrity and confidentiality presents unique challenges. Traditional confidentiality measures, such as noise addition, protect privacy by obscuring model updates but can inadvertently impede the detection of malicious patterns, including backdoor attacks or label poisoning. On the other hand, integrity defenses aim to identify and mitigate these threats but may conflict with confidentiality measures, complicating the design of robust systems that achieve both objectives. This complexity is further amplified by the non-i.i.d. nature of real-world data, where the diversity of benign model updates makes it challenging to distinguish legitimate updates from malicious ones, especially when inaccuracies are introduced by confidentiality defenses. These trade-offs create significant barriers to ensuring security and privacy without compromising global model performance.
Despite the critical importance of these issues, not many previous works target to simultaneously address integrity and confidentiality threats while preserving model performance. Nevertheless, tackling this intertwined challenge is imperative for real-world federated learning applications, where system reliability and user trust depend on effective, balanced defense mechanisms.
@@
Leveraging these properties, we propose Tazza, a novel framework that ensures both confidentiality and integrity in federated learning. Specifically, Tazza implements a weight shuffling module that rearranges neural network parameters according to a predefined shuffling rule, effectively obscuring sensitive information. This shuffling process is designed to be computationally efficient, making it suitable for deployment on mobile and embedded devices. Additionally, Tazza includes a shuffled model validation process, which verifies the integrity of the shuffled models by ensuring consistent outputs, thereby maintaining model performance in ubiquitous computing scenarios.
Nevertheless, properly shuffling model parameters while preserving the desired properties is a challenging task, as it requires a deep understanding of the intricate operations within neural networks. Moreover, while many network architectures exhibit permutation equivariant and invariant properties, the underlying mechanisms that enable these properties vary significantly across different models. For this, we introduce tailored algorithms for representative neural network architectures, providing a systematic approach to better understand and exploit these properties. Using our insights and experience, we demonstrate how these algorithms can effectively harness permutation-based properties, paving the way for practical applications in federated learning.
To implement the aforementioned properties, Tazza employs a novel weight shuffling module, designed to (intentionally) obscure the original model by rearranging its weights while maintaining permutation equivariant and invariant properties. The operations of this module ensure that the shuffled model remains valid, which is maintained by a shuffled validation process for integrity checks, confirming that model outputs remain accurate despite weight shuffling. Finally, we leverage the similarity metrics obtained from the shuffled validation process and cluster the models to distinguish and isolate malicious updates from impacting benign models.
By integrating these components, we present Tazza as a secure, privacy-preserving, and accurate federated learning framework that effectively addresses both integrity and confidentiality challenges, thereby fostering trust and facilitating broader ubiquitous computing among users of mobile and embedded devices. Along with a strong mathematical foundation leveraging permutation properties, we demonstrate its practical efficacy through extensive evaluations on diverse datasets and embedded platforms. Specifically, Tazza maintains high model accuracy while mitigating both integrity and confidentiality attacks, with a computational speedup of up to 6.7$\times$ compared to alternative defense schemes. Our key contributions are summarized as follows:
@@ 2. Background and Related Works
2.1. Security Threats in Federated Learning
We begin by discussing background information on the two core security threats in federated learning: integrity and confidentiality attacks. We present a summary of existing federated learning frameworks and defense mechanisms that target to mitigate the aforementioned attacks in Table 1.
Table 1. Comparison of Tazza with existing federated learning frameworks and defense schemes.
To mitigate integrity attacks, prior research has proposed secure model aggregation techniques leveraging statistical priors or similarity-based approaches (c.f., Table 1). Methods like Median (Xie et al., 2018) and Trimmed-Mean (Yin et al., 2018) aggregation exclude malicious updates by focusing on statistical measures rather than naive averaging, assuming that malicious updates are statistical outliers. Alternatively, similarity-based techniques such as Krum (Blanchard et al., 2017) use Euclidean distance to select updates with minimal deviation across models, while FLTrust (Cao et al., 2020) calculates trust scores based on cosine similarity with a verified, clean model stored at the server. Despite their strengths, these methods often discard valuable information in the updates, leading to suboptimal global model performance.
To counter confidentiality attacks, widely adopted approaches include applying differential privacy (Abadi et al., 2016), which adds controlled noise to model updates, and gradient pruning (Zhu et al., 2019), which suppresses sensitive information in gradients during local training. However, these defense mechanisms often degrade model performance, as they manipulate original updates in ways that hinder convergence and accuracy (Huang et al., 2021). Furthermore, such approaches often necessitate repetitive computations during local training, which not only delay the federated training process but also impose additional computational burdens on local devices (c.f., Table 1).
2.2. Permutation Equivariance/Invariance
@@
Permutation Equivariance refers to the property where the order of a function’s output aligns with the order of its input. Figure 2 (a) illustrates this concept using matrix multiplication, where $C$ represents a column-reversed matrix. Specifically, the figure demonstrates that when the column order of the input matrix $X$ is reversed ($X\rightarrow X^{C}$), the resulting matrix multiplication outputs, $WX$ (top) and $WX^{C}$ (bottom), differ only in the column order, which corresponds to the input’s order.
Formally, if a function $f(\cdot)$ satisfies permutation equivariance, for a given input $X=\{x_{1},x_{2},\dots,x_{N}\}$ and a permutation function $\pi(i):\{1,2,\ldots,N\}\to\{1,2,\ldots,N\}$, where $\pi(i)$ is a bijective mapping, the lemma $\pi(f(X))=f(\pi(X))$ holds. Applying this to the case of matrix multiplication, the function $f(X)=WX$ satisfies permutation equivariance with respect to the column order. Consequently, $WX^{C}$ (i.e., shuffling order of input) is identical to $(WX)^{C}$ (i.e., shuffling order of output).
Permutation Invariance describes a function that produces identical results regardless of the order of its input. For example, a function that computes the average of the elements in a given array is permutation invariant because the arrangement of elements does not influence the outcome. Figure 2 (b) illustrates this concept using the average operation in the context of matrix multiplication, where $n$ represents the number of rows and $R$ denotes a row-permuted version of the input. As shown, despite reordering the rows, the final result remains unchanged (i.e., $\mathbb{E}[X]=\mathbb{E}[X^{R}]$). Without loss of generality, if a function $f(\cdot)$ satisfies permutation invariance, then for a given input $X$ and a permutation function $\pi(i)$, the lemma $f(X)=f(\pi(X))$ holds.
It is important to note that matrix multiplication, in general, is not a permutation invariant operation. Specifically, if the weights used in the multiplication are not uniform (e.g., $\frac{1}{n}$ for all elements, as in the averaging operation shown in Figure 2 (b)), the result will differ depending on the order of the rows. This is because such weights can be interpreted as a weighted sum, being sensitive to the input arrangement, emphasizing the importance of exquisite shuffling operation in a neural network to retain the property.
@@
Figure 3. Visualization of a simple neural network topology and its matrix computations to exemplify the permutation invariance and equivariance properties. (a) Original network architecture and computations. (b) Row-permuted operations, where the input nodes ($x_{1}$, $x_{2}$, $x_{3}$) are rearranged to ($x_{2}$, $x_{3}$, $x_{1}$), demonstrating the permutation invariance property as the network topology and output remain unchanged. (c) Column-permuted operations, where the output nodes are reordered, highlighting the permutation equivariance property, as the network maintains consistent output relationships regardless of the arrangement. Best viewed in color.
In Figure 3, we visualize the topology and matrix computations of a simple neural network to provide deeper insights into its structural properties. Specifically, Figure 3 (a) depicts the original architecture and computations of the neural network, while Figures 3 (b) and (c) illustrate row-permuted and column-permuted operations, emphasizing the permutation invariance and equivariance properties, respectively.
[…]
On the other hand, when the output nodes are rearranged, as in Figure 3 (c), this operation corresponds to a column permutation in the weight matrix. Despite the reordering, the underlying network topology remains unchanged, and the outputs are consistent except for their order. This behavior exemplifies the permutation equivariance property of the neural network, where the outputs reflect the same relationships regardless of the node arrangement.
@@ 3. Feasibility Study and Core Idea
Figure 4. (a) Bottom-5 (top green box) and Top-5 (bottom red box) PSNR samples from the MNIST dataset. (b) PSNR distribution between the original and shuffled images from the MNIST dataset (using 60K images).
Feasibility Study. To evaluate the feasibility of leveraging shuffling methods for preserving data privacy, we measured the Peak Signal-to-Noise Ratio (PSNR) between original and shuffled samples in the MNIST dataset. For shuffling, we randomly rearranged the pixels in the input data following a consistent and fixed order applied uniformly across all samples. Figure 4 presents sample original/shuffled images and the distribution of measured PSNR values between 60,000 original and shuffled images. A high PSNR indicates structural similarity between two images, whereas a low PSNR suggests significant differences. As Figure 4 (a) shows, the shuffled samples do not retain the semantic information of the original images. Moreover, Figure 4 (b) shows that most samples exhibit low PSNR in the range of 6–8 dB. Even for the top-5 PSNR samples, the shuffled images remain semantically uninterpretable to human observers, confirming the effectiveness of shuffling in preserving data privacy.
@@
Next, we trained a 3-layer MLP model for 20 epochs using two dataset configurations: one with original data and the other with shuffled data, as exemplified in Figure 4 (a). For the shuffling process, the test data and training data were shuffled using the same order across all samples. Figure 5 presents the average validation accuracy for 10 different random seeds. As the plots show, the model achieves (nearly) identical accuracy regardless of whether the input data were shuffled, demonstrating the neural network’s semi-permutation invariance property. This result suggests that data shuffling can preserve computation consistency in both training and inference, while effectively addressing user data privacy concerns since the training data is spatially shuffled which hinders the semantic information as shown in Figure 4.
Figure 6. Visualization of sample data illustrating the impact of operation order on the data semantics. (a) Shuffling applied prior to augmentation operations (center crop and horizontal flip), resulting in disrupted spatial coherence and semantically meaningless data. (b) Shuffling applied after the augmentation operations, preserving semantic integrity and highlighting the importance of processing order.
@@
To illustrate this, we present a visualization of samples subjected to shuffling and two commonly used augmentation techniques (i.e., center crop and horizontal flip) applied in different sequences. Figure 6 (a) and (b) depict scenarios where the shuffling operation is performed either prior to or after the augmentation steps, respectively. As demonstrated, applying shuffling prior to cropping or flipping disrupts the spatial coherence of the data, leading to semantically meaningless samples. Conversely, when shuffling is executed following the augmentations, the semantic integrity of the data is preserved. This underscores the critical importance of the operational sequence in maintaining data utility.
@@ 4. Threat Model
Goals and Capability First, an integrity attacker refers to a malicious client that aims to compromise the performance of the global model by injecting harmful information through tampered model updates. Integrity attackers can launch model or data poisoning attacks, manipulating the local training process by altering labels, training data, or model weights. Second, we consider an honest-but-curious server as a confidentiality attacker, which seeks to infer a client’s training data by exploiting the received information. Given that clients transmit their model updates to the server for aggregation, the attacker has access to the updates explicitly shared by the clients. Furthermore, there can be additional shared information, which can be any type of data that the clients are required to transmit for the global model training process (e.g., the amount of training data used).
Assumptions For the integrity attack, we assume that the attacker has no control over its device beyond the local training process. Specifically, the attacker cannot interfere with the weight aggregation, distribution, or client selection processes carried out by the federated learning server, nor can it access information from other benign clients, such as their data distribution, model weights, or updates. Finally, we assume that the proportion of attackers cannot exceed 50% of the total number of clients (Fang et al., 2020), all being common assumptions in prior work. In the case of a confidentiality attacker, the attacker cannot interfere with or access the local operations performed on clients, nor can it eavesdrop on network communications. Lastly, we note that both types of attackers make every effort to achieve their objectives and circumvent defense mechanisms, while the victim systems can operate cooperatively to prevent these attacks or leverage additional security features, for instance, peer-to-peer communication (Wu et al., 2022) and multi-party computation (Zhao et al., 2019). We note that the extensive lines of previous studies have tried to establish such features in practical, scalable, and efficient aspects; thus, we do not consider this as a strong assumption (Kabir et al., 2024).